OpenStudy (meilendurcer):

The following function defines a recursive sequence: f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? A. -2, 8, -26, -80, … B. -2, 8, -26, 80, … C. -2, 12, -44, 180, … D. -2, -12, -44, -180, …

OpenStudy (tkhunny):

Well, the ones that don't have f(1) = 8 seem to be poor candidates.

OpenStudy (meilendurcer):

yeah

OpenStudy (tkhunny):

Calculate f(2) = -4•f(1) - 3•f(0)

OpenStudy (meilendurcer):

uhm... yeah i don't know how to do that

OpenStudy (tkhunny):

Yes you do. What is f(0)? What is f(1)?

OpenStudy (meilendurcer):

oh -2 and 8 sorry went blank

OpenStudy (tkhunny):

f(2) = -4•8 - 3•(-2) = ??

OpenStudy (meilendurcer):

-26?

OpenStudy (tkhunny):

Okay. Now is the time to make a final decision. Find f(3) = -4•f(2) - 3•f(1)

OpenStudy (meilendurcer):

A?

OpenStudy (meilendurcer):

B!!!!! yay!!

OpenStudy (meilendurcer):

I realize this is what you just helped with but it looks more confusing. The first four terms of a sequence are shown below: 7, 4, 1, -2 3 and -3?? Which of the following functions best defines this sequence? f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1

OpenStudy (tkhunny):

What does this mean "3 and -3??" 7-4 = 3 4-1 = 3 1-(-2) = 3 I'm seeing a pattern.

OpenStudy (meilendurcer):

7 - 4 = 3 i did math wrong for 1,-2

OpenStudy (meilendurcer):

subtracting 3?

OpenStudy (meilendurcer):

so basically its the second answer lol that didn't take long

OpenStudy (tkhunny):

That's all there is to it. :-)

OpenStudy (meilendurcer):

Since youre here and seem pretty good with math can you help me with my final exams?

OpenStudy (tkhunny):

You're funny. Have to do exams on your own. Do well, though.

OpenStudy (meilendurcer):

aha thanks. Hopefully. thanks for your help!! ^_^